Supplemental Operations Report – May 12th

Date: 5/12/2024
Name of person filing report: Sergii Iakymov
Reason for Report: Routine
Non-Nominal Systems: Hab wall heater, solar observatory, perseverance parking brake & battery. EVA suit 9, Solar Power inverters slave 1&2, main generator offline

Power system:
Solar: Two malfunctioning batteries replaced. Inverters Slave 1&2 have internal faults, removed from the system and are being send to repairs. With one working inverter we are down to 4.4kW maximum we can use.
SOC Last 24 hours: Max 100%; Min 87%; Avg 96.1%.
VDC Last 24 hours: Max 58.49V; Min 48.69V; Avg 53.46V.
Generator run time: Is out of service. It is not turning on due to Locked rotor shutdown. Service call scheduling is in progress. Oil, oil filter, air filter, spark plugs, frequency garnet replaced on 05/05/2024

Propane Readings:
Station Tank: 75%
Director Tank: 70%
Intern Tank: 78%
Generator Tank: 73%

Hab Static Tank – 540 gallons
GreenHab – 0 gallons
Outpost tank – 400 gallons
Science Dome – 0 gallons
Septic Tanks emptied: No

Sojourner rover used: Yes
Hours: 200.4
Beginning Charge: 100 %
Ending Charge: 100 %
Currently Charging: No
Notes on Rovers: Perseverance is charging only up to about 90%. Charging error appears from time to time. Curiosity rear driver side tire replaced due to cruck. All rovers batteries replaced. All five rovers moved to GreenHab area for charging from the portable generator.

ATV: ATV’s Used: None. Nothing to report.

Hab Car used and why, where: To Hanksville for supplies.
Crew Car used and why, where: To Grand Junction, Price and Hanksville for crew and supplies.
General notes and comments: Hab car ignition has hard time to turn on from time to time.

Summary of Internet: All nominal. Starlink moved to Hab north side.
EVA suits and radios: Suits: Suit 9 has damaged visor and put aside until end of the season.
Comms: Two headsets have been replaced.
Campus wide inspection, if action taken, what and why: Solar Power down to 4.4kW. Main generator out of service. Portable generator is being used to charge rovers.
Summary of Hab Operations: Wall heater is not functioning. Diagnostic is pending. Starlink moved to the Hab north side.
Summary of GreenHab Operations: Shutted down for summer.
Summary of SciDome Operations: Two malfunctioning batteries replaced. Two Inverters with internal fault removed.
Summary of Observatories Operations: Solar observatory telescope diagonal replaced. Lower shutter is decoupled and fastened by a bungee cord. Astronomy laptop has motherboard and battery issues.
Summary of RAM Operations: All nominal
Summary of Outpost Operations: All nominal
Summary of Health and Safety Issues: All nominal

Copyright © The Mars Society. All rights reserved. | Main Site